For many years, Thomson Company manufactured a single product called LEC 40. Then three years ago, the company automated a portion of its plant and at the same time introduced a second product called LEC 90 that has become increasingly popular. The LEC 90 is a more complex product, requiring 0.80 hours of direct labor time per unit to manufacture and extensive machining in the automated portion of the plant. The LEC 40 requires only 0.40 hours of direct labor time per unit and only a small amount of machining. Manufacturing overhead costs are currently assigned to products on the basis of direct labor-hours.

Despite the growing popularity of the company’s new LEC 90, profits have been declining steadily. Management is beginning to believe that there may be a problem with the company’s costing system. Direct material and direct labor costs per unit are as follows:

LEC 40 LEC 90
Direct materials $ 24.00 $ 44.00
Direct labor (0.40 hours and 0.80 hours @ $10.00 per hour) $ 4.00 $ 8.00

Management estimates that the company will incur $744,000 in manufacturing overhead costs during the current year and 60,000 units of the LEC 40 and 20,000 units of the LEC 90 will be produced and sold.

2. Management is considering using activity-based costing to assign manufacturing overhead cost to products. The activity-based costing system would have the following four activity cost pools:

Activity Cost Pool Activity Measure Estimated Overhead Cost
Maintaining parts inventory Number of part types $ 126,000
Processing purchase orders Number of purchase orders 120,000
Quality control Number of tests run 42,000
Machine-related Machine-hours 456,000
$ 744,000
Expected Activity
Activity Measure LEC 40 LEC 90 Total
Number of part types 400 650 1,050
Number of purchase orders 1,400 600 2,000
Number of tests run 700 1,400 2,100
Machine-hours 4,000 2,000 6,000

a. Determine the activity rate for each of the four activity cost pools.

Using the activity rates you computed in part (2), do the following:

a. Determine the per unit amount of manufacturing overhead cost that would be assigned to each product using the activity-based costing system.

b. Compute the unit product cost of each product.
Answer #1

ANS A.

ACTIVITY RATES :                                    Basis                                                      Activity Rate

1. Maintaining Parts Inventory                     No of Parts                   126000/1050 = 120

2 . Processing purchase orders    No of Purchase oders                             120000/2000 = 60

3 . Quality controls                                   No of Test runs                                       42000/2100 = 20

4. Machine related                                   Machine Hours                                       456000/6000 = 76

Activity Rates for Product LEC40

1. Maintaining Parts Inventory                     No of Parts                   75600/400 =   189

2 . Processing purchase orders    No of Purchase oders                             72000/1400 = 51.42

3 . Quality controls                                   No of Test runs                                       25200/700   = 36.00

4. Machine related                                   Machine Hours                                       273600/4000 = 68.40

total = 344.82

Activity Rates for Product LEC90

1. Maintaining Parts Inventory                     No of Parts                   50400/650 = 77.53

2 . Processing purchase orders    No of Purchase oders                             48000/600 = 80

3 . Quality controls                                   No of Test runs                                       16800/1400 = 12

4. Machine related                                   Machine Hours                                       182400/2000 = 91.2

total : 260.73

B COST OF EACH PRODUCT

LEC 40 LEC 90
Direct materials $ 24.00 $ 44.00
Direct labor (0.40 hours and 0.80 hours @ $10.00 per hour) $ 4.00 $ 8.00
Prime cost $ 28.00 $ 52.00
Over Head cost $ 344.82 260.73
Total cost $ 372.82 312.73

Company is manufacturing of two products

1 LEC40

2 LEC 90

LABOUR HOURS FOR PRODUCING THE PRODUCT

FOR,

LEC 40 : 0.4 HOURS

LEC 90 :0.8 HOURS

TOTAL UNITS PRODUCED :

LEC 40 : 60000

LEC 90 : 20000

TOTAL LABOUR HOURS

FOR

LEC 40 : 60000* 0.4: 24000 HRS

LEC90 : 20000*0.8 :16000 HRS

OVER HEAD APPORTION

Activity Cost Pool                         Amount        Basis                 for LEC40             forLEC90

1. Maintaining parts inventory          126000         Labour Hours      75600 50400

2. Processing purchase oders    120000    72000    48000

3 Quality Control    42000                                   25200                 16800

4 Machine Related    456000    273600 182400